OpenStudy (wasiqss):

(4xy^2+6y)dx+(5yx^2+8x)dy=0 solve by using integrating factor

OpenStudy (wasiqss):

sam got clue?

sam (.sam.):

(4xy^2+6y)+(5yx^2+8x)dy/dx=0

OpenStudy (wasiqss):

sam whats in that?

OpenStudy (ash2326):

\[y(4xy+6) dx+x(5xy+8)dy=0\] M=y(4xy+6) N=x(5xy+8) Integrating factor=\(\frac{1}{Mx-Ny}\) Mx=xy(4xy+6) Ny=xy(5xy+8) Integrating factor=\[\frac{1}{xy(4xy+6-(5xy+8))}\] IF=\[\frac{1}{-2xy-x^2y^2}\] Multiply this by our DE , \[\frac{y(4xy+6)}{2xy+x^2y^2} dx+\frac{x(5xy+8)}{2xy+x^2y^2}dy=0\] First term is M and second is N solution \[\int M (y\ treated\ as\ constant)+\int( terms\ of\ N\ not\ containing\ x)dy=C\] ,

OpenStudy (wasiqss):

hold on till i integrate

OpenStudy (wasiqss):

ash having hard time integrating it :/

OpenStudy (ash2326):

Yeah After multiplying the integrating factor we get \[\frac{(4xy+6)y}{xy(xy+2)} dx+\frac{(5xy+8)x}{xy(xy+2)} dy=0\] We'll get now \[(\frac{4y}{xy+2}+\frac{6}{x^2y+2x})dx+(\frac{5x}{2+xy}+\frac{8}{y(xy+2)})dy\] The term with dx we have to treat y as constant and integrate it In the second term we have to integrate only the terms without x, it may seem taht there is no term without x but we can get a term without x \[(\frac{4y}{xy+2}+\frac{6/y}{x^2+2x/y})dx+(\frac{5x}{2+xy}+\frac{4(2+xy-xy)}{y(xy+2)})dy\] \[(\frac{4y}{xy+2}+\frac{6/y}{x^2+2x/y+1/y^2-1/y^2})dx+(\frac{5x}{2+xy}+\frac{8}{y}-\frac{4xy}{y(xy+2)})dy\] Now Let's integrate \[4y \ln(xy+2) \times 1/y+ \int \frac{6/y}{(x+1/y)^2-(1/y)^2}dx+ \int \frac{8}{y} dy=c\] Wasiqq do the integration of middle term, I don't remember its formula and your DE is solve

OpenStudy (ash2326):

Can you do it now? Did you understand?

OpenStudy (wasiqss):

im tryin to understand it

OpenStudy (wasiqss):

hey jus temme how to integrate 1/(yx^2+2x)

OpenStudy (wasiqss):

lalaly i want an easy way out to do it

OpenStudy (lalaly):

ok i need 10 min...

OpenStudy (wasiqss):

ok time alloted :P

OpenStudy (wasiqss):

its 20 min now :P

OpenStudy (wasiqss):

@lalaly

OpenStudy (jamesj):

What is written above is right. There isn't an "easy" way. This is the way.

OpenStudy (wasiqss):

james im having difficulty integrating it :(

OpenStudy (wasiqss):

@JamesJ

OpenStudy (jamesj):

In that penultimate expression \[ \int \frac{6/y}{(x + 1/y)^2 - (1/y)^2} dx \] treat all of the y and 1/y as constants. This will appear less intimidating if you write a = 1/y and change variables to x' = x + 1/y = x + a. Now as dx' = dx, we have \[ \int \frac{6a}{x'^2 - a^2} dx' \]

OpenStudy (wasiqss):

@JamesJ final thing how to integrate 1/(yx^2+2x)

OpenStudy (jamesj):

withe respect to what? y or x?

OpenStudy (wasiqss):

with respect to x

OpenStudy (wasiqss):

taking Y constant

OpenStudy (wasiqss):

@JamesJ

OpenStudy (jamesj):

then complete the square. yx^2 + 2x = y ( x^2 + (2/y)x ) = y [ ( x^2 + (2/y)x + (1/y)^2 ) - (1/y)^2 ] = y [ ( x + (1/y) )^2 - (1/y)^2 ]

OpenStudy (wasiqss):

james look at my other ques